# Simultaneously search for multi-target Galactic binary gravitational waves in reduced parameter space with LMPSO-CV

### v1

@@ VI.1 False source fraction and SNR
Figure 16: The SNR ratio is calculated for confirmed sources ($R$geq 0.8) in comparison to their corresponding injection sources. The term “unremove degeneracy noise” signifies the SNR before the $-\mathcal{F}$-statistic analysis, whereas “remove degeneracy noise” indicates the SNR after the $-\mathcal{F}$-statistic analysis.
The distribution of reported sources in different SNR and $R$ ranges is illustrated in Fig. 17. Each cell on SNR has a width of 2, and on $R$, it is 0.01. The highest density of the distribution is in the range where $R$ spans from 0.98 to 0.99, and SNR from 13 to 15, as well as in the range where $R$ ranges from 0.97 to 0.98, and SNR from 11 to 13. Some reported sources have SNR $>$ 15, which may be influenced by instrument noise and overlapped signals.
Figure 17: The distribution of reported sources in different SNR and $R$ ranges.
The FAS is also sensitive to the SNR of the sources. We put the number of these sources in different conditions and the associated FAS in Tab. 1. The highest FAS occurs when the SNR range is between 7-9, and the $R$ threshold is set at 0.9. Due to the lower SNR of these signals, they deviate more from the parameter position of the binary, exhibiting higher sensitivity to instrument noise. The lowest FAS occurs when SNR $>$ 13 and the $R$ threshold is 0.5. The higher SNR and lower $R$ threshold result in a FAS of $6.2\%$. In conclusion, the FAS increases with the decrease of SNR and the increase of $R$ threshold.
Correlation coefficient ($R$)
Signal to noise ratio (SNR)
Reported sources ($N_{r}$)
Confirmed sources ($N_{c}$)
False source fraction (FAS)
$R$geq 0.5
7-9
2554
1342
47.5%
9-11
2381
1772
25.6%
11-13
1873
1591
15.1%
$>$13
1980
1857
6.2%
$R$geq 0.8
7-9
2554
654
74.4%
9-11
2381
1177
50.6%
11-13
1873
1158
38.2%
$>$13
1980
1533
22.6%
$R$geq 0.9
7-9
2554
370
85.5%
9-11
2381
787
66.9%
11-13
1873
859
54.1%
$>$13
1980
1150
41.9%
Table 1: The number of sources and the false source fraction in different thresholds of $R$ and ranges of SNR.
VI.2 Parameter errors
When different thresholds for $R$ are applied, the reported sources exhibit varying parameter errors. We categorize the reported sources into three groups based on $R$: $R$geq 0.9, 0.8 $\leq$ $R$ $<$ 0.9, and 0.5 $\leq$ $R$ $<$ 0.8, and investigate the magnitude of parameter errors for each category in Fig. 18. For sources with $R$geq 0.9, characterized by high correlation coefficients, the parameter errors are relatively lower. Their average errors in $f$, $\dot{f}$, $\beta$, and $\lambda$ are $3.11\times 10^{-9}$ Hz, $9.42\times 10^{-17}$ $\text{Hz}^{2}$, 0.07 rad, and 0.03 rad, respectively. In the case of sources with 0.8 $\leq$ $R$ $<$ 0.9, the average errors in $f$, $\dot{f}$, $\beta$, and $\lambda$ are $3.85\times 10^{-9}$ Hz, $1.10\times 10^{-16}$ $\text{Hz}^{2}$, 0.13 rad, and 0.08 rad, respectively. For sources with 0.5 $\leq$ $R$ $<$ 0.8, characterized by low correlation coefficients, the parameter errors are higher. Their average errors in $f$, $\dot{f}$, $\beta$, and $\lambda$ are $2.16\times 10^{-8}$ Hz, $1.22\times 10^{-16}$ $\text{Hz}^{2}$, 0.31 rad, and 0.28 rad, respectively. The secondary peaks on either side of the main peak of $\Delta f$ in Fig. 18 for sources with $0.5\leq$ $R$ $<$ 0.8 suggest that some sources, whose $f$ values deviate from those of the real binary by an amount greater than the minimum frequency resolution $\mathrm{d}f=1.59\times 10^{-8}$ Hz for the two-year detection, have been searched in the degeneracy noise position. For sources with $0.8\leq$ $R$ $<$ 0.9, the rate of change in the average errors of $f$, $\dot{f}$, $\beta$, and $\lambda$ parameters, relative to sources with $R$ $\geq 0.9$, are 23.8%, 16.8%, 85.7%, and 166.7%, respectively. Therefore, the main difference between relatively accurate sources and accurate sources lies in the accuracy of the sky position. For sources with $0.5\leq$ $R$ $<$ 0.8, the rate of change in the average errors of $f$, $\dot{f}$, $\beta$, and $\lambda$ parameters, relative to sources with $0.8\leq$ $R$ $<$ 0.9, are 461.0%, 10.9%, 138.5%, and 250.0%, respectively. Except for $\dot{f}$, the errors in the other three parameters have shown a significant increase.
@@ VI.4 Discussion
The numbers of confirmed sources in different frequency bins are depicted in Fig. 19. The injection sources refer to signals with 7 $<$ optimal SNR $<$ 15 in the LDC1-4 data. With a low correlation coefficient threshold of $R$ = 0.5, it becomes feasible to detect a substantial portion or approximately half of the injection sources in medium ($1\times 10^{-3}$ Hz $<f<4\times 10^{-3}$ Hz) and low ($f<1\times 10^{-3}$ Hz) frequency ranges. However, in the frequency bins with the highest signal density ($\sim 2\times 10^{-3}$ Hz), a notable number of sources remain undetected ($>50\%$). This observation might be associated with our established stop search criteria or potential interference among the signals themselves. At higher frequencies ($f>4\times 10^{-3}$ Hz), our method demonstrates poor performance. This is attributed to the widening frequency distribution of degeneracy noise as the signal frequency increases. Consequently, it becomes easier to identify the “secondary peaks” (degeneracy noise) produced by the signal rather than the “primary peak” (real signal). To enhance performance at high frequencies, allocating additional computing resources is advisable, and adjustments to our stop search criteria may be necessary. However, by constraining the $\dot{f}$ range and conducting a logarithmic search without altering other settings, we successfully identify the sources at higher frequencies. In Fig. 19, the red line serves as the cutoff at $4\times 10^{-3}$ Hz. On the right side of the line, we utilize the outcomes after removing individual degeneracy noise, restricting the Galactic Latitude, and removing overlapped degeneracy noise in the second search (no cross-validation). On the left side of the line, we employ the confirmed sources obtained through complete $-\mathcal{F}$-statistic-analysis.
@@
Figure 19: The number of sources in different frequency bins. The injection sources denoting those with 7 $<$ optimal SNR $<$ 15 in LDC1-4 data. The other three colored lines illustrate the count of confirmed sources under different $R$ thresholds to the left of the red vertical line. The red line serves as the cutoff at $4\times 10^{-3}$ Hz. On the right side of the line, we utilize the outcomes after removing individual degeneracy noise, restricting Galactic Latitude, and removing overlapped degeneracy noise in the second search.
@@
In Fig. 20, we present the Power Spectral Density (PSD) of various datasets. The term “LDC1-4 data” refers to raw data from the LISA Data Challenge, while “instrument noise” represents simulated noise for the LISA mission. “Residual SNR15” denotes LDC1-4 residual data, where injection sources with an optimal SNR $\geq$ 15 have been removed—representing the dataset on which our method operates. The term “Residual confirmed” signifies data from which all confirmed sources ($R\geq 0.8$) have been subtracted from “Residual SNR15” (as depicted in “remove degeneracy noise” in Fig. 16). We update the amplitude $\mathcal{A}$ of each source in proportion to the square root of the change in their $\mathcal{F}$-statistic values before and after the removal of degeneracy noise. Finally, “Residual SNR7” corresponds to LDC1-4 residual data with injection sources removed based on an optimal SNR $\geq$ 7, making it an ideal target for our search methodology. There are instances where some sources are not detected, or the reported sources in the results lack accuracy ($R$ $<$ 0.8). This phenomenon is attributed to the limitations of the current create voids and $-\mathcal{F}$-statistic-analysis methods, particularly in resolving signals with very close parameters. For instance, in the case of the signal presented in Fig. 9, where the frequency difference between the two signals is less than $1\times\mathrm{d}f$, treating them as individual signals becomes challenging. In such scenarios, effectively eliminating degeneracy noise originating from both signals becomes intricate. Even in situations where two or three signals overlap, despite the frequency difference exceeding $1\times\mathrm{d}f$ (as illustrated in Fig. 7(d) and Fig. 10(d)), inconsistencies between the $\mathcal{F}$-statistic peak and the binary position add complexity to the subsequent process of removing degeneracy noise. These complexities contribute to an increase in the FAS. To overcome these challenges, future approaches, such as modeling and simultaneously searching for multiple overlapped sources (global fit) or leveraging machine learning to analyze the distribution of $\mathcal{F}$-statistic values directly in the local maxima, represent promising avenues.
@@ VII Conclusion
We introduce the LMPSO-CV approach designed for the simultaneous detection of Galactic binary gravitational waves using a space-based GW detector. In dealing with the numerous local maxima of the $\mathcal{F}$-statistic, the majority of which are degeneracy noise, we employ a $-\mathcal{F}$-statistic-analysis to effectively remove both degeneracy and instrument noise. We depict the flow chart of our method in Fig. 14. Notably, our approach mitigates the challenges associated with inaccurate subtraction contamination, particularly in the context of resolving low SNR signals (e.g., SNR $<$ 15), a common difficulty in traditional iterative-subtraction methods. The $-\mathcal{F}$-statistic-analysis comprises four key steps: 1) eliminating individual signal degeneracy noise, 2) conducting a second search with cross-validation, 3) restricting the Galactic Latitude, and 4) removing overlapped signals degeneracy noise. Following the completion of these steps, our method identifies 8995 signals with a false source fraction (FAS) of $47.7\%$ when applied to the processed LDC1-4 data, which involves subtracting the waveforms of injection sources with optimal SNR $>$ 15. Alternatively, selecting signals with a change in $\mathcal{F}$-statistic value of less than 10% before and after the $-\mathcal{F}$-statistic-analysis yields 3463 reported sources with a FAS $=26.9\%$. In instances where two or more signals exhibit similarity in their parameters, particularly concerning frequency, the occurrence of overlapped signals is inevitable. The situation becomes intricate when the frequency separation between two signals with comparable Signal-to-Noise Ratio (SNR) is less than five times the minimum frequency resolution achievable during a two-year observation ($\Delta f<5\times\mathrm{d}f$). Under such circumstances, the $\mathcal{F}$-statistic peaks associated with these overlapped signals become complex, introducing challenges in the process of signal resolution. This intricacy adds to the overall challenge of effectively reducing the false source fraction.
@@ Acknowledgments
P. Gao thanks S. D. Mohanty, X. Zhang, J. Ming, S. Zhao and Y. Guo for their invaluable comments and suggestions. Our Local Maxima Particle Swarm Optimization (LMPSO) algorithm is developed based on the PSO script in GBSIEVER method [38] (this PSO code is actually taken from the public repository https://github.com/mohanty-sd/SDMBIGDAT19.git). This work is supported by National Key R&D Program of China (2020YFC2201400), and by the NSFC (No. 11920101003, No. 11922303, No. 12021003 and No. 12005016). X. Fan. is supported by the Fundamental Research Funds for the Central Universities (No.2042022kf1182). Z. Cao was supported by “the Interdiscipline Research Funds of Beijing Normal University” and CAS Project for Young Scientists in Basic Research YSBR-006. We also gratefully appreciate the Wuhan Supercomputing Center for the generous computing resources to support our research efforts.

### v3

@@ VI.1 Detection rate and SNR
Figure 15: The SNR-ratio is calculated for confirmed sources ($R_{t}=0.8$) in comparison to their corresponding injection sources. The term “unremove degeneracy noise” signifies the SNR before the $-\mathcal{F}$-statistic analysis, whereas “remove degeneracy noise” indicates the SNR after the $-\mathcal{F}$-statistic analysis.
VI.2 Compared with other methods
There have been recent developments in the search for LDC1-4 data [38, 44, 43]. As our focus primarily lies on sources with SNR $<15$, direct comparison with their findings is not feasible. However, using Fig. 16, we can roughly assess the effectiveness of these methods. The figure displays the number of injection sources in the LDC1-4 data across different SNR thresholds and various observation durations, comparing results from four methods: Result 1 from Table 2 of Ref. [43], Result 2 from Table IV of Ref. [44], Result 3 from Table I of Ref. [38], and our own results from this paper. In Fig. 16, Result 1 slightly exceeds the number of injection sources with SNR $\geq$ 15 for both 6-month (6m) and 1-year (1y) detection durations, indicating strong capability in identifying sources above this threshold. We anticipate it will also yield slightly more sources than the injection sources with SNR $\geq$ 15 during a two-year (2y) detection period. Result 2 shows a similar trend to the injection sources with SNR $\geq$ 15; however, due to a higher correlation coefficient threshold ($R\geq 0.9$), it is slightly lower than that count. Results 2 and 3 are nearly identical for a two-year observation period. Our result is derived from the deduction of all injection sources with SNR $\geq$ 15. Therefore, in Fig. 16, we use the count of injection sources with SNR $\geq$ 15 as a reference point, and the total number is 15,076. Since our result includes only sources with SNR $<$ 15, which are more susceptible to noise and denser—leading to increased interference (see Sec. III.4)—its detection rate is naturally lower compared to the other three methods.
By referring to Table I in Ref. [38], we can make a rough assessment of the effectiveness of our method, too. They present the main outcomes of their method applied to the LDC1-4 data, categorizing sources’ frequencies into three distinct regions. Similarly, we categorize our result into the same three regions, as shown in Tab. 2. The Ref. [38] has 2778 reported sources with $f<3$ mHz and SNR $<$ 25. Since there are approximately 3000 injection sources within the SNR range of 15 $\sim$ 25 and $f<3$ mHz, it appears that Ref. [38] might not have extensively searched for sources with $f<3$ mHz and SNR $<15$. Conversely, our method identifies 5206 reported sources within the range, exhibiting varying performance depending on the chosen $R_{t}$. Hence, our approach may perform better in the low frequency and low SNR range. A similar comparison can be drawn for the region 3 $43\text{ mHz}<f<$ 4 mHz. For $f>4$ mHz, there are approximately 1000 injection sources with SNR $<15$, where our method does not perform as effectively as that of Ref. [38]. We provide an explanation for this discrepancy in Sec. VI.5. Overall, our approach demonstrates a comparative advantage in low frequency and low SNR regions.
Figure 16: This figure illustrates the number of injection sources (with the ordinate in base-10 logarithm) at different SNR thresholds across various observing durations. The SNR is calculated based on instrument noise. It also compares the number of sources identified by four different search methods: Result 1 is from Ref. [43], Result 2 is from Ref. [44], Result 3 is from Ref. [38], and our own result are included as well. Additionally, we highlight the number of injection sources for different SNR thresholds specifically for a two-year observing period.
Frequency $<3$ mHz $3\sim 4$ mHz $>4$ mHz All
C $N_{r}$ 5206 1297 5 6508
DR ($R_{t}=0.5$) 79.0% 95.9% 100.0% 82.4%
DR ($R_{t}=0.8$) 57.1% 87.5% 80.0% 63.2%
DR ($R_{t}=0.9$) 38.2% 78.3% 80.0% 46.3%
Injection sources’ SNR 3 $\sim$ 15
Table 2: The number and detection rate of reported sources (C) in Fig. 14, when we take different $R_{t}$ and ranges of frequency. We utilize injection sources with SNR ranging from 3 to 15 to compute the correlation coefficient, ensuring that our confirmed sources fall within this range and do not correspond to injection sources with higher SNR.
Indeed, the number of confirmed sources searched in Refs. [38, 44] and the injection sources eliminated by us to acquire the LDC1-4 residual data are comparable. According to our hypothesis, we can first employ the Refs. [38, 44, 43] or other methods to identify sources with high SNR or high frequency and subsequently utilize our method to detect those with low SNR and low frequency in the residual. Given the higher density of low SNR sources, the overlap among them will exceed that of high SNR sources (as shown in Sec. III.4). Iterative subtraction search method risks subtracting “secondary peaks”, which will lead to the energy loss of real signal and potentially affect the position of the “primary peak”. In contrast, our method entails identifying all peaks without subtraction, followed by retaining the “primary peak” and discarding the “secondary peaks”. This strategy avoids the effect of inaccurate subtraction contamination.
Our pipeline can also be applied to high SNR sources (e.g., SNR $\geq$ 15), but the significant amount of degeneracy noise originating from high SNR sources will impose greater demands on computing resources. Therefore, we believe it is preferable to initially employ the Ref. [38] or other methods to process high SNR sources and subsequently utilize our pipeline to search for low SNR sources in the residual data. Our pipeline is primarily designed for scenarios where numerous sources with similar SNR coexist, making it particularly effective for detecting low SNR and low-frequency sources.
VI.3 Parameter errors
When we taking different $R_{t}$, the reported sources exhibit varying parameter errors. We categorize the reported sources into three groups based on $R$: $R$geq 0.9, 0.8 $\leq$ $R$ $<$ 0.9, and 0.5 $\leq$ $R$ $<$ 0.8, and investigate the magnitude of parameter errors for each category in Fig. 17. For sources with $R$geq 0.9, characterized by high correlation coefficients, the parameter errors are relatively lower. Their average errors in $f$, $\dot{f}$, $\beta$, and $\lambda$ are $3.11\times 10^{-9}$ Hz, $9.45\times 10^{-17}$ $\text{Hz}^{2}$, 0.07 rad, and 0.03 rad, respectively. In the case of sources with 0.8 $\leq$ $R$ $<$ 0.9, the average errors in $f$, $\dot{f}$, $\beta$, and $\lambda$ are $3.90\times 10^{-9}$ Hz, $1.10\times 10^{-16}$ $\text{Hz}^{2}$, 0.13 rad, and 0.09 rad, respectively. For sources with 0.5 $\leq$ $R$ $<$ 0.8, characterized by low correlation coefficients, the parameter errors are higher. Their average errors in $f$, $\dot{f}$, $\beta$, and $\lambda$ are $1.77\times 10^{-8}$ Hz, $1.18\times 10^{-16}$ $\text{Hz}^{2}$, 0.28 rad, and 0.28 rad, respectively. The secondary peaks on either side of the main peak of $\Delta f$ in Fig. 17 for sources with $0.5\leq$ $R$ $<$ 0.8 suggest that some sources, whose $f$ values deviate from those of the real binary by an amount greater than the minimum frequency resolution $\mathrm{d}f=1.59\times 10^{-8}$ Hz for the two-year detection, have been searched in the degeneracy noise position.
For sources with $0.8\leq$ $R$ $<$ 0.9, the rate of change in the average errors of $f$, $\dot{f}$, $\beta$, and $\lambda$ parameters, relative to sources with $R$ $\geq 0.9$, are 25.4%, 16.4%, 85.7%, and 200.0%, respectively. Therefore, the main difference between relatively accurate sources and accurate sources lies in the accuracy of the sky position. For sources with $0.5\leq$ $R$ $<$ 0.8, the rate of change in the average errors of $f$, $\dot{f}$, $\beta$, and $\lambda$ parameters, relative to sources with $0.8\leq$ $R$ $<$ 0.9, are 353.8%, 7.3%, 115.4%, and 211.1%, respectively. Except for $\dot{f}$, the errors in the other three parameters have shown a significant increase.
@@ VI.5 Discussion
The numbers of confirmed sources in different frequency bins are depicted in Fig. 18. The injection sources refer to signals with 7 $<$ SNR $<$ 15 in the LDC1-4 data. With a low correlation coefficient threshold of $R_{t}$ = 0.5, it becomes feasible to detect a substantial portion or approximately half of the injection sources in medium ($1\times 10^{-3}$ Hz $<f<4\times 10^{-3}$ Hz) and low ($f<1\times 10^{-3}$ Hz) frequency ranges. However, in the frequency bins with the highest signal density ($\sim 2\times 10^{-3}$ Hz), a notable number of sources remain undetected ($>50\%$). This observation might be associated with our established stop search criteria or potential interference among the signals themselves. At higher frequencies ($f>4\times 10^{-3}$ Hz), our method demonstrates poor performance. This is attributed to the widening frequency distribution of degeneracy noise as the signal frequency increases. Consequently, it becomes easier to identify the “secondary peaks” (degeneracy noise) produced by the signal rather than the “primary peak” (real signal). To enhance performance at high frequencies, allocating additional computing resources is advisable, and adjustments to our stop search criteria may be necessary. However, by constraining the $\dot{f}$ range and conducting a logarithmic search without altering other settings, we successfully identify the sources at higher frequencies. In Fig. 18, the red line serves as the cutoff at $4\times 10^{-3}$ Hz. On the right side of the line, we utilize the outcomes after removing individual degeneracy noise, restricting the Galactic Latitude, and removing overlapping degeneracy noise in the second search (no cross-validation). On the left side of the line, we employ the confirmed sources obtained through complete $-\mathcal{F}$-statistic-analysis.
@@
Figure 18: The number of sources in different frequency bins. The injection sources denoting those with 7 $<$ SNR $<$ 15 in LDC1-4 data. The other three colored lines illustrate the count of confirmed sources under different $R$ ranges to the left of the red vertical line. The red line serves as the cutoff at $4\times 10^{-3}$ Hz. On the right side of the line, we utilize the outcomes after removing individual degeneracy noise, restricting Galactic Latitude, and removing overlapping degeneracy noise in the second search.
@@
In Fig. 19, we present the Power Spectral Density (PSD) of various datasets. The term “LDC1-4 data” refers to raw data from the LISA Data Challenge, while “instrument noise” represents simulated noise for the LISA mission. “Residual SNR15” denotes LDC1-4 residual data, where injection sources with SNR $\geq$ 15 have been removed—representing the dataset on which our method operates. The term “Residual confirmed” signifies data from which all confirmed sources ($R_{t}=0.8$) have been subtracted from “Residual SNR15” (as depicted in “remove degeneracy noise” in Fig. 15). We update the amplitude $\mathcal{A}$ of each source in proportion to the square root of the difference in their $\mathcal{F}$-statistic values before and after the removal of degeneracy noise. Finally, “Residual SNR7” corresponds to LDC1-4 residual data with injection sources removed based on SNR $\geq$ 7, making it an ideal target for our search methodology. There are instances where some sources are not detected, or the reported sources in the results lack accuracy ($R$ $<$ 0.8). This phenomenon is attributed to the limitations of the current create voids and $-\mathcal{F}$-statistic-analysis methods, particularly in resolving signals with very close parameters. For instance, when the frequency difference between two signals is less than $1\times\mathrm{d}f$, it becomes challenging to treat them as individual signals. In such scenarios, effectively eliminating degeneracy noise originating from both signals becomes intricate. Even in situations where two or three signals overlap, despite the frequency difference exceeding $1\times\mathrm{d}f$ (as illustrated in Fig. 7(d) and Fig. 9(d)), inconsistencies between the $\mathcal{F}$-statistic peak and the binary position add complexity to the subsequent process of removing degeneracy noise. These complexities contribute to the decrease in the DR. To overcome these challenges, other approaches, such as multi-source fitting or leveraging machine learning to analyze the distribution of $\mathcal{F}$-statistic values directly in the local maxima, represent promising avenues.
@@ VII Conclusion
We introduce the LMPSO-CV approach designed for the simultaneous detection of Galactic binary gravitational waves using a space-based GW detector. In dealing with the numerous local maxima of the $\mathcal{F}$-statistic, the majority of which are degeneracy noise, we employ a $-\mathcal{F}$-statistic-analysis to effectively remove both degeneracy and instrument noise. We depict the flow chart of our method in Fig. 13. Notably, our approach mitigates the challenges associated with inaccurate subtraction contamination, particularly in the context of resolving low SNR signals (e.g., SNR $<$ 15), a common difficulty in traditional iterative-subtraction methods. The $-\mathcal{F}$-statistic-analysis comprises four key steps: 1) eliminating individual signal degeneracy noise, 2) conducting a second search with cross-validation, 3) restricting the Galactic Latitude, and 4) restricting the difference in $\mathcal{F}$-statistic and removing overlapping signals degeneracy noise. Following the completion of these steps, our method identifies 6508 signals with DR $=63.2\%$ ($R_{t}=0.8$) when applied to the LDC1-4 residual data, which involves subtracting the waveforms of injection sources with SNR $\geq$ 15.
Our approach offers an advantage in low SNR (e.g., SNR $<$ 15) and low-frequency ranges (e.g., $f<4\times 10^{-3}$ Hz). Once the high SNR (e.g., SNR $\geq$ 15) and high frequency (e.g., $f\geq 4\times 10^{-3}$ Hz) sources have been subtracted, it is worthwhile to continue the search using our method in the residual.
In instances where two or more signals exhibit similarity in their parameters, particularly concerning frequency, the occurrence of overlapping signals is inevitable. The situation becomes intricate when the frequency separation between two signals with comparable Signal-to-Noise Ratio (SNR) is less than five times the minimum frequency resolution achievable during a two-year observation ($\Delta f<5\times\mathrm{d}f$). Under such circumstances, the $\mathcal{F}$-statistic peaks associated with these overlapping signals become complex, introducing challenges in the process of signal resolution. This intricacy adds to the overall challenge of effectively improving the detection rate.
@@ Acknowledgments
P. Gao thanks S. D. Mohanty, X. Zhang, J. Ming, S. Zhao and Y. Guo for their invaluable comments and suggestions. Our Local Maxima Particle Swarm Optimization (LMPSO) algorithm is developed based on the PSO script in GBSIEVER method [38] (this PSO code is actually taken from the public repository https://github.com/mohanty-sd/SDMBIGDAT19.git). This work is supported by National Key R&D Program of China (2020YFC2201400), and by the NSFC (No. 11920101003, No. 11922303, No. 12021003 and No. 12005016). X. Fan. is supported by the Fundamental Research Funds for the Central Universities (No.2042022kf1182). Z. Cao was supported by “the Interdiscipline Research Funds of Beijing Normal University” and CAS Project for Young Scientists in Basic Research YSBR-006. We gratefully appreciate the Wuhan Supercomputing Center for the generous computing resources to support our research efforts. We also thank LISA Consortium’s LDC working group for curating and supporting the simulated data used in this study.